# Efficient Analysis of Latent Spaces in Heterogeneous Networks

### v2

@@ 4 Theory for estimating latent spaces
For a matrix $X\in\mathbb{R}^{n\times m}$, let $\mathcal{G}(X)=X^{\top}X/n\in\mathbb{R}^{m\times m}$ denote the Gram matrix of $X$ standardized by its row number and let $\sigma_{\operatorname{min}}(X)$ denote the minimum singular value of $X$. Assume that there exist positive constants $M_{1}$, $M_{2}$, and $M_{3}$ such that
@@
Condition 2 (i) implies that $\ell_{2}$-norms of the latent vector $z_{i}^{\star}$ and $w_{t,i}^{\star}$ are bounded over all $i$ uniformly. It constrains all the parameters to be in a bounded space, which is a prevalent mild assumption in high-dimensional network analysis (Ma et al.,, 2020; Zheng and Tang,, 2022). Condition 2 (ii) and (iii) impose constraints on the scaled Gram matrices of the latent factors. They imply that the columns of $[Z^{\star},W_{t}^{\star}]$ (for any $t$) and $[Z^{\star},W_{t}^{\star},W_{s}^{\star}]$ (for one pair of $\{t,s\}$) are linearly independent, ensuring that the true model parameters are identifiable by Proposition 1. For Condition 2 (ii), note that
@@
MacDonald et al., (2022) studies a similar model with generalized inner product (Rubin-Delanchy et al.,, 2022). Their theoretical analysis focuses on edge-wise normal distribution with $\nu(x)=x$, and their imposed conditions imply that as $n\to\infty$,
@@
Assume the log-partition function $\nu:\mathbb{R}\to\mathbb{R}$ in the model (1) satisfies the following conditions.
@@
    $\nu(\theta)$ is cubic differentiable with a continuous third derivative;
@@
    there exist constants $\kappa_{1}$, $\kappa_{2}$, and $\kappa_{3}$ such that $0<\kappa_{1}\leqslant\nu^{\prime\prime}(\theta)\leqslant\kappa_{2}$ and $|\nu^{\prime\prime\prime}(\theta)|\leqslant\kappa_{3}$ for any $\theta\in[-2M_{1}^{2},2M_{1}^{2}]$, where $M_{1}$ is given in Condition 2;
  • (iii)
    there exists a constant $L>0$ such that for any $1\leqslant i\leqslant j\leqslant n$ and $1\leqslant t\leqslant T$, $\mathbb{E}\{|A_{t,ij}-\nu^{\prime}(\Theta_{t,ij}^{\star})|^{l}\}\leqslant\nu^{\prime\prime}(\Theta_{t,ij}^{\star})L^{l-2}l!/2$ for any integer $l\geqslant 2$, where $l!=l(l-1)\cdots 1$ represents the factorial of $l$.
Condition 3 (i) and (ii) assume the smoothness and convexity of the log-partition function $\nu$, which is commonly used in likelihood-based analysis. Condition 3 (iii) assumes a Bernstein moment condition for all edges $A_{t,ij}$, which provides convenience to use concentration inequalities in theoretical proofs. Condition 3 can be satisfied by most exponential families of distributions. For instance, $\nu(x)={x^{2}}/{2}$ for normal distribution, $\nu(x)=\log(1+\exp(x))$ for Bernoulli distribution, and $\nu(x)=\exp(x)$ for Poisson distribution (Efron,, 2022).
@@
Assume Conditions 13. Let $(\mathring{Z},\mathring{W})$ be the estimators obtained through Algorithm 1 with $\mathring{Y}_{t}$ in Theorem A.2 as initialization. For any constant $\varepsilon>0$, there exists a constant $C_{\varepsilon}>0$ such that when $n/\log^{d_{\operatorname{max}}}(T)$ with $d_{\operatorname{max}}=\max_{1\leqslant t\leqslant T}d_{t}$ is sufficiently large,
@@
Theorem 1 suggests that with high probability, the overall estimation error rate of $\operatorname{dist}^{2}(\mathring{Z},Z^{\star})$ and $\operatorname{dist}^{2}(\mathring{W}_{t},W_{t}^{\star})$ are of the order of $O(1)$ up to logarithmic factors. This is the expected oracle rate for $\Theta_{t}^{\star}$ as discussed in Section 2.3.
@@
Theorem 1 only provides an overall upper bound on the estimation error rates for shared and heterogeneous parameters simultaneously. Establishing the desired efficient error rates for each group of parameters separately is challenging, as discussed in Section 2.3. To facilitate the theoretical analysis, we make an additional assumption on the regularity of the heterogeneous parameters.
@@
We compare Proposition 4 with Proposition 3 to gain insights into Condition 4. First, (a) in Proposition 4 specifies an upper bound on the singular values of $W_{t}^{\star\top}W_{t}^{\star}/n$, which is compatible with the lower bound (a) in Proposition 3 when $T$ is sufficiently large. Second, (b) in Proposition 4 specifies an upper bound on the singular values of $W_{t}^{\star\top}W_{s}^{\star}/n$, which differs from (b) in Proposition 3, as $M_{5}$ depends on constants $(M_{1},M_{2})$ in Condition 2. Condition 4 is primarily a technical condition needed to facilitate the control of certain high-order error terms in the analysis. Intuitively, it ensures that the column spaces of $W_{t}^{\star}$ and $W_{s}^{\star}$ for any pair $t\neq s$ are sufficiently different, making it easier to disentangle their respective estimation errors. We point out that the constraint on $M_{5}$ may not limit the practical use of our method. Numerical studies in Section 5 demonstrate its robust performance even in the extreme case where $W_{t}^{\star}$’s are identical. This empirical evidence suggests that the constraint on $M_{5}$ could potentially be relaxed under certain scenarios, which, however, is not pursued here to preserve the clarity and coherence of our presentation.
@@
Assume Conditions 14. Let $(\hat{Z},\hat{W})$ be the refined estimators from Algorithm 2 with $(\mathring{Z},\mathring{W})$ in Theorem 1 as initialization and the adjustment in Remark 5. For any constant $\varepsilon>0$, there exists a constant $C_{\varepsilon}>0$ such that when $n/\log^{\varsigma}(T)$ with $\varsigma=\max\{d_{\operatorname{max}},8\}$ is sufficiently large,
@@
When $1<T\lesssim n$, (13)$=O(1/T)$ ), showing that the semiparametric oracle rate for the shared factors $Z^{\star}$ discussed in Section 2.3 can be achieved. In this case, $\hat{Z}$ and $\hat{W}_{t}$ can achieve the oracle estimation error rate for estimating $Z^{\star}$ and $W_{t}^{\star}$ up to logarithmic factors. On the other hand, when $T\gg n$, (13)$=O(1/n)$ ), which is referred to as a sub-optimal rate. Intuitively, when $T$ is too large, the degree of heterogeneity becomes larger, and the estimation problem is more challenging. In addition, the aggregated squared estimation error of $W_{t}^{\star}$ is of the order of $O_{p}(1)$ up to logarithmic factors. This is also consistent with the oracle rate of $W_{t}^{\star}=[w_{t,1}^{\star},\ldots,w_{t,n}^{\star}]^{\top}$, $O_{p}(n\times 1/n)=O_{p}(1)$, following Section 2.3.
@@
As Table 1 suggests, our derived error rates in Theorem 2 not only align with the best rates reported in the compared works but also advance the literature by accommodating the most flexible range of link functions. While He et al., (2023) reported similar error rates for both shared and individual parameters, our work distinguishes itself by addressing unique and previously insurmountable challenges highlighted in Section 2.3. We develop a comprehensive set of analytical innovations including effective shared space identification validated by Theorem 1, and a novel joint one-step construction framework addressing the multiplex singularity issue of the efficient information matrix in Remark 2. These breakthroughs collectively establish a solid foundation for comprehensive analysis of heterogeneous networks, deepening our understanding of efficiency gains beyond existing statistical paradigms.
@@ 5 Simulation studies
In Section 4, we have pointed out that the proposed estimators do not require the columns of $Z^{\star}$ and $W_{t}^{\star}$ to be nearly orthogonal, characterized through $\mathcal{G}([Z^{\star},W^{\star}])$. To demonstrate that, we next generate true parameters with $k=k_{t}$ for $t=1,\ldots,T$ and the scaled Gram matrix $\mathcal{G}([Z^{\star},W^{\star}])=\Omega\otimes\mathrm{I}_{k}/(2\sqrt{k})$, where $1/(2\sqrt{k})$ is a normalization factor, and $\Omega=(\omega_{i,l})\in\mathbb{R}^{(1+T)\times(1+T)}$ has diagonal entries equal to one. This Gram structure implies that columns of $Z^{\star}$ are orthogonal, columns of $W_{t}^{\star}$ are orthogonal, and for $1\leqslant t,s\leqslant T$, $\angle(Z_{j}^{\star},W_{t,j}^{\star})=\cos^{-1}(\omega_{1,1+t})$ and $(W_{t,j}^{\star},W_{s,j}^{\star})=\cos^{-1}(\omega_{1+t,1+s}),$ where $Z_{j}^{\star}$ and $W_{t,j}^{\star}$ represent the $j$-columns of $Z^{\star}$ and $W^{\star}_{t}$, respectively, and (,$\angle$($\cdot$,$\cdot$)∠ ( ⋅ , ⋅ ), represents the angle between two input column vectors. When $\omega_{1,1+t}$ is larger, the angle between $Z_{j}^{\star}$ and $W_{t,j}^{\star}$ is smaller, suggesting less orthogonality; similar interpretation applies to $\omega_{1+t,1+s}$. We next consider three cases (A)–(C) with $\Omega$ set as follows:
@@
where $\mathbf{0}$ represents all-zero matrices whose dimensions are set such that $\Omega$ is of size $(1+T)\times(1+T)$, $\mathbf{1}_{T}$ represents a $T$-dimensional all-one column vector, $\Sigma_{T}(\rho)$ denotes a $T\times T$ matrix with diagonal elements equal to $1$ and off-diagonal elements equal to $\rho$, i.e., with a compound symmetry structure. Under Case (A), all columns of $[Z^{\star},W^{\star}]$ are orthogonal. Under Case (B), $\angle(Z_{j}^{\star},W_{t,j}^{\star})=\cos^{-1}(\phi)$ and $\angle(W_{t,j}^{\star},W_{s,j}^{\star})=\cos^{-1}(\rho)$ for $1\leqslant t\neq s\leqslant T$. In simulations, we take $\phi=0.1$ and $\rho=0.3$, corresponding to an intermediate level of non-orthogonality. Under Case (C), the first $1+T_{o}$ columns of $[Z^{\star},W^{\star}]$ are orthogonal with the other columns, whereas $\angle(W_{t,j}^{\star},W_{s,j}^{\star})=\cos^{-1}(1)=0$, i.e., $W_{t,j}^{\star}=W_{s,j}^{\star}$ for $T_{o}+1\leqslant t,s\leqslant T$. In simulations, we take $T_{o}=4$, corresponding to an extreme scenario allowing part of the heterogeneous factors to overlap and thus are highly non-orthogonal. To randomly generate true parameters with a given scaled Gram matrix denoted as $\mathcal{G}^{\star}$, we first generate $\tilde{Z}$ and $\tilde{W}_{t}$ with rows independently sampled from a $k$-dimensional standard normal distribution restricted on the set $\{x\in\mathbb{R}^{k}:\|x\|_{2}^{2}\leqslant k\}$, and then set $[Z^{\star},W^{\star}]=[\tilde{Z},\tilde{W}]\tilde{\mathcal{G}}^{-1/2}(\mathcal{G}^{\star})^{1/2}$, where $\tilde{\mathcal{G}}=\mathcal{G}([\tilde{Z},\tilde{W}])$. Given true parameters in each case, we simulate data under three distributions Bernoulli, Gaussian, and Poisson in (1), corresponding to edge values that are binary, continuous, and event counts, respectively.
In each setting, we evaluate the performance of the proposed Algorithms 1 and 2, referred to as SS-Hunting and SS-Refinement, respectively. The two methods MultiNeSS and MultiNeSS+ in MacDonald et al., (2022) are also compared as our models are similar. Since their codes are not available under Poisson distribution, their errors are not presented under Poisson. We next focus on illustrating the order of the semiparametric error rate of estimating $Z^{\star}$ with respect to the number of networks $T$, highlighting the improvement over using a single network. The error rates for estimating $W_{t}^{\star}$’s are not improved over $T$, which is consistent with our expected oracle rates in Section 2.3, and detailed results are deferred to Section G of the Supplementary Material due to the limitation of space. For the ease of visualization, we vary $T\in\{5,10,20,40,80\}$ while fixing $n=400$ and $k_{t}=k=2$ for $t=1,\ldots,T$. For a fair comparison with MacDonald et al., (2022), we present error $\|\widehat{ZZ^{\top}}-Z^{\star}Z^{\star\top}\|_{\mathrm{F}}^{2}/n$, which is equivalent to $\operatorname{dist}^{2}(\hat{Z},Z^{\star})$ up to constants multiplied (Tu et al.,, 2016).
@@
Figures 13 present empirical estimation errors of the four estimators under Cases (A)–(C), respectively. In Figure 1, all the errors are inverse proportional to $T$, achieving the oracle error rate discussed in Section 2.3. Among the four methods, SS-Refinement and MultiNeSS+ achieve the smallest error under different distributions. In Figure 2, SS-Refinement achieves the smallest error across three distributions, while SS-Hunting performs similarly under Gaussian and Poisson distributions but not Bernoulli distribution. Different from Figure 1, MultiNeSS and MultiNeSS+ no longer exhibit inverse proportional to $T$ error rates. Under Case (C), errors of MultiNeSS and MultiNeSS+ explode and thus are not presented in Figure 3. Unlike Figures 1 and 2, SS-Hunting does not achieve inverse proportional to $T$ errors, whereas SS-Refinement still achieves that.
@@ 6 Data analysis
We analyze a multiplex-network dataset of lawyers (Lazega,, 2001) to demonstrate the insights provided by examining shared and individual latent spaces. This dataset contains three types of connection relationships: coworker, friendship, and advice, between seventy-one lawyers in a Northeastern US corporate law firm. Within each type of network, we consider binary and undirected edges between lawyers indicating whether connections of each type exist between them. Since there is no ground truth of latent spaces in practice, we compare the estimated latent embeddings with observed node-wise features to gain insights into the results given by the proposed method and other existing methods.
As network edges are binary, we adopt Bernoulli distribution in the model (1) and estimate latent dimensions by Algorithm B.1, which gives $\hat{k}_{1}=6$, $\hat{k}_{2}=4$, $\hat{k}_{3}=1$, and $\hat{k}=2$. Then we estimate latent embedding vectors under estimated latent dimensions by SS-Refinement. As a comparison, we also apply MultiNeSS+ in MacDonald et al., (2022) and MASE in Arroyo et al., (2021) with the dimension of the shared space set to be two.
(a) SS-Refinement
@@
(a) SS-Refinement
@@
(a) SS-Refinement
@@
We first illustrate the shared latent embeddings estimated by the three methods. Figure 4 (a) and (b) present the leading two principal component scores of the shared latent component $Z$ estimated by SS-Refinement and MultiNeSS+, respectively. Figure 4 (c) displays latent embeddings estimated by MASE that are rotated to approximate (a) and (b) for the ease of comparison. In Figure 4, each point is colored according to the office location of the corresponding lawyer. The patterns of points in Figure 4 (c) slightly differ from those in (a) and (b), which could be due to the difference between models fitted. Nevertheless, the results of all three methods show that the shared latent embeddings are correlated with the office locations of lawyers. This suggests that office locations could play a shared role in forming multiple types of connections between lawyers.
We next compare the individual latent embeddings $W_{t}$’s estimated by SS-Refinement and MultiNeSS+, where we point out that MASE does not directly give comparable estimates and thus is not included. Figures 5 and 6 present the estimated individual latent component after projected to its leading two principal components for the coworker and friendship networks, respectively. In Figure 5, points are colored according to lawyers’ practice. The results show that latent vectors estimated by both two methods demonstrate association with lawyars’ practice, whereas the separation between litigation and corporate lawyers appears to be more obvious in estimates given by SS-Refinement. This discrepancy might be attributed to the fact that MultiNeSS+ tends to encourage the column spaces of $Z$ and $W_{t}$’s to be orthogonal, which is not required in our approach and thus leads to distinct estimates in practice. In Figure 6, points are colored based on the lawyers’ status. Similarly to Figure 5, latent vectors estimated by both two methods exhibit correlation with a common covariate, lawyers’ status, even though the patterns of points are not exactly the same.

### v3

@@ 4 Theory for estimating latent spaces
For a matrix $X\in\mathbb{R}^{n\times m}$, let $\mathcal{G}(X)=X^{\top}X/n\in\mathbb{R}^{m\times m}$ denote the Gram matrix of $X$ standardized by its row number, and let $\sigma_{\operatorname{min}}(X)$ denote the minimum singular value of $X$. Assume that there exist positive constants $M_{1}$, $M_{2}$, and $M_{3}$ such that
@@
Condition 2 (i) implies that $\ell_{2}$-norms of the latent vector $z_{i}^{\star}$ and $w_{t,i}^{\star}$ are bounded over all $i$ uniformly. It constrains all the parameters to be in a bounded space, which is a prevalent mild assumption in high-dimensional network analysis (Ma et al.,, 2020; Zheng and Tang,, 2022). Condition 2 (ii) and (iii) impose constraints on the scaled Gram matrices of the latent factors. They imply that the columns of $Z^{\star},W_{t}^{\star}]$ (for any $t$) and $Z^{\star},W_{t}^{\star},W_{s}^{\star}]$ (for one pair of $\{t,s\}$) are linearly independent, ensuring that the true model parameters are identifiable by Proposition 1. For Condition 2 (ii), note that
@@
MacDonald et al., (2022) studies a similar model with generalized inner product (Rubin-Delanchy et al.,, 2022). Their theoretical analysis focuses on edge-wise normal distribution with $\mu(\theta)=\theta$, and their imposed conditions imply that as $n\to\infty$,
@@
Let $\mathcal{X}=\{x\in\mathbb{R}:p(\,x\mid\theta\,)>0\}$ denote the support of $p(\,\cdot\mid\theta\,)$. Assume $l(\theta;x)$ in (5) belongs to the natural exponential family and satisfies the following conditions.
@@
    For any fixed $x\in\mathcal{X}$, $l(\theta;x)$ is three times differentiable with respect to $\theta$, with its first to third derivatives with respect to $\theta$ denoted by $l^{\prime}(\theta;x)$, $l^{\prime\prime}(\theta;x)$, and $l^{\prime\prime\prime}(\theta;x)$, respectively. Moreover, there exist positive constants $\kappa_{1}$, $\kappa_{2}$, and $\kappa_{3}$ such that $\kappa_{1}\leqslant-l^{\prime\prime}(\theta;x)\leqslant\kappa_{2}$ and $|l^{\prime\prime\prime}(\theta;x)|\leqslant\kappa_{3}$ for any $x\in\mathcal{X}$ and $|\theta|\leqslant 2M_{1}^{2}$, where $M_{1}$ is given in Condition 2.
@@
    There exists a constant $L>0$ such that $\mathbb{E}|l^{\prime}(\Theta_{t,ij}^{\star};A_{t,ij})|^{m}\leqslant\mathrm{var}\{l^{\prime}(\Theta_{t,ij}^{\star};A_{t,ij})\}L^{m-2}m!/2$ for any $1\leqslant i\leqslant j\leqslant n$, $1\leqslant t\leqslant T$, and any integer $m\geqslant 2$, where $m!=m(m-1)\cdots 1$ represents the factorial of $m$.
Condition 3 can be naturally satisfied by most distributions under natural exponential families. For instance, $l(\theta;x)\propto\theta x-\log(1+\exp(\theta))$ for Bernoulli distribution, and $l(\theta;x)\propto\theta x-\exp(\theta)$ for Poisson distribution (Efron,, 2022). The assumption of natural exponential family is adopted here primarily for the ease of presentation and understanding. We show that generalizations under more general distributions can be established following similar arguments in Section I of the Supplementary Material.
@@
Assume Conditions 13. Let $(\mathring{Z},\mathring{W})$ be the estimators obtained through Algorithm 1 with $\mathring{Y}_{t}$ in Theorem A.2 as initialization. For any constant $\varepsilon>0$, there exist positive constants $c_{\varepsilon}$ and $C_{\varepsilon}$ such that when $\log^{d_{\operatorname{max}}+2}(nT)/n\leqslant c_{\varepsilon}$ with $d_{\operatorname{max}}=\max_{1\leqslant t\leqslant T}d_{t}$,
@@
Theorem 1 suggests that with high probability, the overall estimation error rate of $\operatorname{dist}^{2}(\mathring{Z},Z^{\star})$ and $\operatorname{dist}^{2}(\mathring{W}_{t},W_{t}^{\star})$ are of the order of $O(1)$ up to logarithmic factors. This aligns with the expected oracle rate for $\Theta_{t}^{\star}=Z^{\star}Z^{\star\top}+W_{t}^{\star}W_{t}^{\star\top}$ as discussed in Section 2.3.
@@
Theorem 1 only provides an upper bound on the estimation error rates for shared and heterogeneous parameters simultaneously. Establishing the desired efficiency for each group of parameters separately is challenging, as discussed in Section 2.3. To facilitate the theoretical analysis, we make an additional assumption on the heterogeneous parameters.
@@
We compare Proposition 4 with Proposition 3 to gain insights into Condition 4. First, (a) in Proposition 4 specifies an upper bound on the singular values of $W_{t}^{\star\top}W_{t}^{\star}/n$, which is compatible with the lower bound (a) in Proposition 3 when $T$ is sufficiently large. Second, (b) in Proposition 4 specifies an upper bound on the singular values of $W_{t}^{\star\top}W_{s}^{\star}/n$, which differs from (b) in Proposition 3, as $M_{5}$ depends on constants $(M_{1},M_{2})$ in Condition 2. Condition 4 is primarily a technical condition needed to facilitate the control of certain high-order error terms in the analysis. Intuitively, it ensures that the column spaces of $W_{t}^{\star}$ and $W_{s}^{\star}$ for most pairs $t\neq s$ are sufficiently different, making it easier to disentangle their respective estimation errors. We point out that the constraint on $M_{5}$ may not limit the practical use of our method. Numerical studies in Section 5 demonstrate its robust performance even in the extreme case where $W_{t}^{\star}$’s are identical. This empirical evidence suggests that the constraint on $M_{5}$ could potentially be relaxed under certain scenarios, which, however, is not pursued here to preserve the clarity and coherence of our presentation.
@@
[…]
@@
[…]
@@
As Table 1 suggests, Theorem 2 not only aligns with the best rates reported in the compared works but also advances the literature by accommodating the most flexible range of link functions. While He et al., (2025) reported similar error rates for both shared and individual parameters, our work distinguishes itself by addressing unique and previously insurmountable challenges highlighted in Section 2.3. We develop a comprehensive set of analytical innovations including effective shared space identification validated by Theorem 1, and a novel joint one-step construction framework addressing the multiplex singularity issue of the efficient information matrix in Remark 3. These breakthroughs collectively establish a solid foundation for comprehensive analysis of heterogeneous networks, deepening our understanding of efficiency gains beyond existing statistical paradigms.
@@ 5 Simulation studies
In Section 4, we highlight that the proposed estimators do not require the columns of $Z^{\star}$ and $W_{t}^{\star}$ to be nearly orthogonal, characterized through $\mathcal{G}([Z^{\star},W^{\star}])$. To demonstrate that, we next generate true parameters satisfying $\mathcal{G}([Z^{\star},W^{\star}])=\Omega\otimes\mathrm{I}_{k}/(2\sqrt{k})$, where $\Omega=(\omega_{i,l})\in\mathbb{R}^{(1+T)\times(1+T)}$ with $\omega_{i,i}=1$, and $1/(2\sqrt{k})$ is a normalization scalar. This Gram structure implies that columns of $Z^{\star}$ are orthogonal, columns of $W_{t}^{\star}$ are orthogonal, and $\cos(Z_{j}^{\star},W_{t,j}^{\star})=\omega_{1,1+t}$ and $\cos(W_{t,j}^{\star},W_{s,j}^{\star})=\omega_{1+t,1+s}$, where $\cos(a,b)=\langle a,b\rangle/(\|a\|_{2}\|b\|_{2})$. When $\omega_{1,1+t}$ is larger, the two vectors $Z_{j}^{\star}$ and $W_{t,j}^{\star}$ are less orthogonal; similar interpretation applies to $\omega_{1+t,1+s}$. We next consider three cases (A)–(C) with $\Omega$ set as follows:
@@
[…]
In each setting, we evaluate the performance of the proposed Algorithms 1 and 2, referred to as SS-Hunting and SS-Refinement, respectively. The two methods MultiNeSS and MultiNeSS+ in MacDonald et al., (2022) are compared as our models are similar. Their errors are not presented under Poisson distribution as codes are unavailable. We next focus on illustrating the error rate of estimating $Z^{\star}$ with respect to $T$, highlighting the improvement over using a single network. The error rates for estimating $W_{t}^{\star}$’s are not improved over $T$, which is consistent with our expected oracle rates in Section 2.3, and detailed results are deferred to Section G of the Supplementary Material. For the ease of visualization, we vary $T\in\{5,10,20,40,80\}$ while fixing $n=400$ and $k_{t}=k=2$ for $t=1,\ldots,T$. For a fair comparison with MacDonald et al., (2022), we present error $\|\widehat{ZZ^{\top}}-Z^{\star}Z^{\star\top}\|_{\mathrm{F}}^{2}/n$, equivalent to $\operatorname{dist}^{2}(\hat{Z},Z^{\star})$ up to constants multiplied (Tu et al.,, 2016).
@@
Figures 13 present empirical estimation errors of the four estimators under Cases (A)–(C), respectively. In Figure 1, all the errors are inverse proportional to $T$, achieving the oracle error rate discussed in Section 2.3, and SS-Refinement and MultiNeSS+ achieve the smallest errors. In Figure 2, SS-Refinement achieves the smallest error across three distributions, while SS-Hunting performs similarly under Gaussian and Poisson distributions but not Bernoulli distribution. Different from Figure 1, errors of MultiNeSS and MultiNeSS+ are no longer inverse proportional to $T$. Under Case (C), errors of MultiNeSS and MultiNeSS+ explode and are not presented in Figure 3 for visual clarity. More details of their results are given in Section G.3 of the Supplementary Material. Unlike in Figures 1 and 2, errors of SS-Hunting are not inverse proportional to $T$, whereas SS-Refinement still achieves that, demonstrating efficiency gain from the likelihood refinement.
@@ 6 Data analysis
We analyze a multiple-network dataset of lawyers (Lazega,, 2001), which contains three types of connection relationships: coworker, friendship, and advice, between seventy-one lawyers in a US corporate law firm. Within each type of network, we consider binary and undirected edges between lawyers indicating whether connections of each type exist between them. Since there is no ground truth of latent spaces in practice, we compare the estimated latent embeddings with observed node-wise features for interpretation.
As network edges are binary, we adopt Bernoulli distribution with logistic link and estimate latent dimensions by Algorithm B.1, which gives $\hat{k}_{1}=6$, $\hat{k}_{2}=4$, $\hat{k}_{3}=1$, and $\hat{k}=2$. Then we estimate latent embedding vectors under estimated latent dimensions by SS-Refinement. As a comparison, we also apply MultiNeSS+ in MacDonald et al., (2022) and MASE in Arroyo et al., (2021) with the dimension of the shared space set to be two.
(a) $S$-Refinement
@@
We first illustrate the shared latent embeddings estimated by the three methods. Figure 4 (a) and (b) present the leading two principal component scores of the shared latent component $Z$ estimated by SS-Refinement and MultiNeSS+, respectively. Figure 4 (c) displays latent embeddings estimated by MASE that are rotated to approximate (a) and (b) for the ease of comparison. In Figure 4, each point is colored according to the office location of the corresponding lawyer. The results of all three methods show that the shared latent embeddings are correlated with the office locations of lawyers. This suggests that office locations could play a shared role in forming multiple types of connections between lawyers.
(a) $S$-Refinement
@@
We next compare the individual latent embeddings $W_{t}$’s estimated by SS-Refinement and MultiNeSS+, where we point out that MASE does not directly give comparable estimates and thus is not included. Figures 5 and 6 present the estimated individual latent component after projected to its leading two principal components for the coworker and friendship networks, respectively; see more details in Section H.2 of the Supplementary Material. In Figure 5, points are colored according to lawyers’ practice. The results show that latent vectors estimated by both two methods demonstrate association with lawyars’ practice, whereas the separation between litigation and corporate lawyers appears to be more obvious in estimates given by SS-Refinement. This discrepancy might be attributed to the fact that MultiNeSS+ tends to encourage the column spaces of $Z$ and $W_{t}$’s to be orthogonal, which is not required in our approach and thus leads to distinct estimates in practice. In Figure 6, points are colored based on the lawyers’ status. Similarly to Figure 5, latent vectors estimated by both two methods exhibit correlation with a common covariate, lawyers’ status, even though the patterns of points are not exactly the same.
(a) $S$-Refinement